# thanks: M.C. and D.P. contributed equally to this work

### v1

@@ A.2 Asymptotic behaviour near the horizon
where we kept only terms scaling as $\sim f^{0},g^{0}$. It is easy to see that in the cases of the Simpson-Visser and Peltola-Kunstatter space-times, for which $f=g$, the second term in square brackets vanishes. With some effort, one can show that this is true for the D’Ambrosio-Rovelli metric as well. This then leaves us with the following equation for $Y_{s}$:
@@ A.2.2 Peltola-Kunstatter
For the Peltola-Kunstatter space-time the functions $f(r)$ and $g(r)$ are given by the following:
@@ A.2.3 D’Ambrosio-Rovelli
For the D’Ambrosio’Rovelli space-time the functions $f(r)$ and $g(r)$ are given by the following:
@@
with $r_{H}^{\text{DR}}=\sqrt{4M^{2}-\ell^{2}}$. From Eq. (59) it follows that $\text{K}=2\sqrt{2}M^{3/2}(2M-\ell)$, which cancels once again with the $g^{\prime}\sqrt{f/g}$ term in Eq. (47). Analogously to the previous case, one can therefore conclude that the asymptotic solutions of the radial Teukolsky equation for $R_{s}$ in the near horizon region is indeed given by Eq. (55), thereby confirming the asymptotic scaling quoted in Eq. (20).

### v2

@@ A.2 Asymptotic behaviour near the horizon
where we kept only terms scaling as $\sim f^{0},g^{0}$. It is easy to see that in the cases of the SV and PK space-times, for which $f=g$, the second term in square brackets vanishes. With some effort, one can show that this is true for the DR metric as well. This then leaves us with the following equation for $Y_{s}$:
@@ A.2.2 Peltola-Kunstatter
For the PK space-time the functions $f(r)$ and $g(r)$ are given by the following:
@@ A.2.3 D’Ambrosio-Rovelli
For the DR space-time the functions $f(r)$ and $g(r)$ are given by the following:
@@
with $r_{H}^{\text{DR}}=\sqrt{4M^{2}-\ell^{2}}$. From Eq. (59) it follows that $\text{K}=2\sqrt{2}M^{3/2}(2M-\ell)$, which cancels once again with the $g^{\prime}\sqrt{f/g}$ term in Eq. (47). Analogously to the previous case, one can therefore conclude that the asymptotic solutions to the radial Teukolsky equation for $R_{s}$ in the near horizon region are indeed given by Eq. (55), thereby confirming the asymptotic scaling quoted in Eq. (20).